OpenStudy (anonymous):

solve the diff. eq.: du/dt = e^-2u.

OpenStudy (watchmath):

Try to sevarate the varibale \(e^{2u}du=dt\) Now integrate!

OpenStudy (anonymous):

du/dt = e^(-2u) e^(2u) du = dt Integrate both sides, don't forget the constant.

OpenStudy (anonymous):

i came up with u = (ln(2t + c)) / 2... sound right?